# Finite-Precision Arithmetic Transceiver for Massive MIMO Systems

### v3

@@ I Introduction
  • Extension of finite-precision arithmetic to the multi-user scenario. For the general multi-user scenario, to eliminate inter-user interference, we utilize ZF detection for MU-SIMO systems and ZF precoding [^2] In fact, the achievable sum rate can be slightly improved at a low SNR by using regularized ZF, i.e., MMSE. However, this approach necessitates knowledge of the noise variances for all users at the transmitter, mandating additional feedback. Thus, in this paper, we focus on ZF and MRC/MRT. for MU-MISO systems, respectively. Note that ZF is engaged in matrix inversion, rendering the analysis of its rounding error challenging. Hence, we transform ZF into the least squares (LS) problem and propose the normal equations (NE) method-based precoding and detection. Then, the rounding error bounds and the lower bounds of the achievable sum rate are derived. These bounds are explicit functions of BS antennas, SNR, channel conditions, arithmetic precision, and the number of users. Moreover, asymptotic analyses are presented by our derived bounds.
@@ III Single-User Scenario with Finite-precision Arithmetic
(a) Uplink SIMO with finite-precision arithmetic
(b) Downlink MISO with finite-precision arithmetic
@@ III-B MISO Using MRT with Finite-Precision Arithmetic
Compared between Lemma 3 and Lemma 2, it can be observed that rounding errors are fixed and independent of the number of antennas $M$ in MISO systems. This is because of its only involvement in vector-scalar products, i.e., a series of finite-precision scalar-scalar products, for MISO systems.
@@ IV-A MU-SIMO Using ZF Detection with Finite-Precision Arithmetic
One traditional approach for solving the LS problem is the NE method [34, Alg. 5.3.1]. In this context, we present the NE method-based ZF detection, as outlined in Algorithm 1. Notably, in the massive MIMO systems where $M\gg K$, the NE method requires only half flops of the QR factorization approach. Furthermore, we can derive the rounding error bound in the lemma below.
@@
Compute the matrix-vector products
@@
Compute the matrix-matrix products
@@
Compute the Cholesky factorization
@@
$c_{1}^{\mathrm{u}}$ $=2K\left(\gamma_{2M}+\gamma_{6K+1}/\left(1-2K\gamma_{2K+1}\right)\right).$
And the received signal for the $k$th user is given by
@@ IV-B MU-MISO Using ZF Precoding with Finite-Precision Arithmetic
Compute the matrix-matrix products
@@
Compute the Cholesky factorization
@@
Solve the triangular equations
@@ V-B2 Computational Cost Analysis
$\mathcal{C}_{S}^{l}\left(b\right)=b-1,\mathcal{C}_{S}^{h}\left(\frac{2n}{b}\right)=G\left(\frac{2n}{b}-1\right),\mathcal{C}_{M}^{l}\left(b\right)=b,$ (70)
@@ VI-B3 Impact of $\rho$
Notably, the derived lower bound is loose. The bound itself is weaker than it might have been because of the necessity of restricting the mass of detail to a reasonable level and because of the limitations imposed by expressing the errors in terms of matrix norms [20, 41]. Rounding error analysis primarily serves the purpose of uncovering potential instabilities in algorithms, aiming to refine them based on the intrinsic properties obtained [41]. Furthermore, following the work of [28], we can set a small value $\lambda$ to fit the real rounding error but it is not a lower bound. For example, as illustrated in Fig. 4a, the curve of ‘Fitting bound $\mathtt{fp16},(\lambda=0.5)$’ approaches closely to the curves of ‘Simulated’ in the small and medium $M$ but exceeds it in the large $M$.
@@ VI-D Mixed-Precision Arithmetic Architecture
Finally, we present an illustrative example involving both SIMO and MU-SIMO systems to highlight the influence of the proposed mixed-precision arithmetic architecture. Fig. 7 demonstrates the rates of SIMO and MU-SIMO systems using the proposed mixed-precision arithmetic architecture versus $M$. The results indicate a significant enhancement in communication performance compared with only low-precision arithmetic, approaching levels similar to those achieved with pure full-precision arithmetic.
Comparing Fig. 7a and Fig. 7b, we can find that a larger performance gap between mixed-precision and full-precision in MU-SIMO systems since ZF detection in MU-SIMO systems involves more matrix computations than SIMO systems. Despite this, mixed-precision architecture has significantly compensated for the impact of low-precision arithmetic in MU-SIMO systems.
@@ Appendix C Proof of Proposition 3
$\mathbb{E}\left\{\left\|\mathbf{r}\right\|_{2}^{2}\right\}$ $=\mathbb{E}\left\{\left\|\mathbf{A}^{H}\mathbf{y}\right\|_{2}^{2}\right\}=\mathbb{E}\left\{\left\|\sqrt{\rho}\mathbf{x}+\mathbf{A}^{H}\mathbf{n}\right\|_{2}^{2}\right\}$ (99)

### v5

@@ I Introduction
  • Extension of finite-precision arithmetic to the multi-user scenario. For the general multi-user scenario, to eliminate inter-user interference, we utilize ZF detection for MU-SIMO systems and ZF precoding for MU-MISO systems, respectively. Note that ZF is engaged in matrix inversion, rendering the analysis of its rounding error challenging. Hence, we transform ZF into the least squares (LS) problem and propose the normal equations (NE) method-based precoding and detection. Then, the rounding error bounds and the lower bounds of the achievable sum rate are derived. These bounds are explicit functions of BS antennas, SNR, channel conditions, arithmetic precision, and the number of users. Moreover, asymptotic analyses are presented by our derived bounds.
@@ III-A SIMO Using MRC with Finite-Precision Arithmetic
(a) Uplink SIMO with finite-precision arithmetic
(b) Downlink MISO with finite-precision arithmetic
@@ III-B MISO Using MRT with Finite-Precision Arithmetic
Compared between Lemma 3 and Lemma 2, it can be observed that rounding errors are fixed and independent of the number of antennas $M$ in MISO systems. This is because of its only involvement in vector-scalar products, i.e., a series of finite-precision scalar-scalar products, for MISO systems. Moreover, the rounding errors in MISO systems are not influenced by noise, as MRT is implemented at the transmitter.
@@ IV-A MU-SIMO Using ZF Detection with Finite-Precision Arithmetic
One traditional approach for solving the LS problem is the NE method [^4] Compared to the iterative method, the NE method is more convenient for rounding error analysis because it relies on standard algorithms such as Cholesky factorization, matrix-matrix multiplication, and matrix-vector multiplication. [34, Alg. 5.3.1]. In this context, we present the NE method-based ZF detection, as outlined in Algorithm 1. Notably, in the massive MIMO systems where $M\gg K$, the NE method requires only half flops of the QR factorization approach. Furthermore, we can derive the rounding error bound in the lemma below.
@@
1 Compute the matrix-vector products
@@
2Compute the matrix-matrix products
@@
3Compute the Cholesky factorization
@@
$c_{1}^{\mathrm{u}}$ $=2K\left(\gamma_{2M}+\gamma_{6K+1}/\left(1-2K\gamma_{2K+1}\right)\right),$
where $K$ is the number of users. And the received signal for the $k$th user is given by
@@ IV-B MU-MISO Using ZF Precoding with Finite-Precision Arithmetic
1 Compute the matrix-matrix products
@@
2Compute the Cholesky factorization
@@
3Solve the triangular equations
@@ V-B2 Computational Cost Analysis
$\small\mathcal{C}_{S}^{l}\left(b\right)=b-1,\mathcal{C}_{S}^{h}\left(\frac{2n}{b}\right)=G\left(\frac{2n}{b}-1\right),\mathcal{C}_{M}^{l}\left(b\right)=Gb,$ (70)
@@ VI-B3 Impact of $\rho$
Notably, the derived lower bound is loose but shows the worst-case communication performance. The bound itself is weaker than it might have been because of the necessity of restricting the mass of detail to a reasonable level and because of the limitations imposed by expressing the errors in terms of matrix norms [20, 41]. Rounding error analysis primarily serves the purpose of uncovering potential instabilities in algorithms, aiming to refine them based on the intrinsic properties obtained [41]. Furthermore, following the work of [28], we can set a small value $\lambda$ to fit the real rounding error but it is not a lower bound. For example, as illustrated in Fig. 4a, the curve of ‘Fitting bound $\mathtt{fp16},(\lambda=0.5)$’ approaches closely to the curves of ‘Simulated’ in the small and medium $M$ but exceeds it in the large $M$.
@@ VI-D Mixed-Precision Arithmetic Architecture
Figure 8: Rate of the MU-MISO systems versus $M$ with mixed-precision arithmetic architecture, $K=4$, $\rho=10\mathrm{dB}$ and $b=32$.
Finally, we demonstrate the superiority of the proposed mixed-precision arithmetic architecture under various conditions, including different values of $M$ and $K$, high SNR levels, computational cost, imperfect channel state information (CSI), and other advanced transceivers.
VI-D1 Impact of $M$
We present an illustrative example involving SIMO, MU-SIMO, and MU-MISO systems with ZF detection/precoding to highlight the influence of the proposed mixed-precision arithmetic architecture. Fig. 7 and Fig. 8 demonstrate the rates of SIMO, MU-SIMO, and MU-MISO systems using the proposed mixed-precision arithmetic architecture versus $M$. The results indicate a significant enhancement in communication performance compared with only low-precision arithmetic, approaching levels similar to those achieved with pure full-precision arithmetic. Comparing Fig. 7a and Fig. 7b, we can find that a larger performance gap between mixed-precision and full-precision in MU-SIMO systems since ZF detection in MU-SIMO systems involves more matrix computations than SIMO systems. Despite this, mixed-precision architecture has significantly compensated for the impact of low-precision arithmetic in MU-SIMO systems.
Figure 9: Rate of MU-SIMO systems versus $K$ with mixed-precision arithmetic architecture, $M=128$, $\rho=10\mathrm{dB}$ and $b=8$.
VI-D2 Impact of $K$
Fig. 9 compares the impact of different finite-precision arithmetic on the communication performance of MU-SIMO systems with varying numbers of users. Lower-precision arithmetic is more sensitive to a larger number of users, leading to a greater performance gap compared to full-precision arithmetic. This sensitivity arises because channels with more users have a higher condition number [42], resulting in increased rounding errors (see (48)). Furthermore, the proposed mixed-precision arithmetic architecture mitigates the impact of low-precision arithmetic under different channel conditions, as shown in Fig. 9.
Figure 10: Rate of MU-SIMO systems versus $M$ with mixed-precision arithmetic architecture, $K=4$, $\rho=10,30\mathrm{dB}$ and $b=8$.
VI-D3 Impact of high SNR $\rho$
As the ZF method demonstrates nearly optimal capacity at high SNR levels, we further present the rate of MU-SIMO systems versus $M$ with mixed-precision arithmetic architecture at $\rho=30$ dB as an example in Fig. 10. The results indicate that the proposed mixed-precision arithmetic remains efficient at high SNR levels.
@@ VI-D3 Impact of high SNR $\rho$
Figure 12: Rates of SIMO and MU-SIMO systems with mixed-precision arithmetic architecture and imperfect CSI. ’P’ means perfect CSI, and ’IP’ means imperfect CSI.
VI-D4 Computational Cost
@@
(a) Rate of MU-SIMO systems using MMSE detection versus $M$ with mixed-precision arithmetic architecture, $K=4$, $\rho=10\mathrm{dB}$ and $b=32$.
(b) Rate of MU-MISO systems using WMMSE versus $M$ with mixed-precision arithmetic architecture, $K=16$, $\rho=10\mathrm{dB}$, the convergence threshold $\epsilon=10^{-3}$ and $b=32$.
(c) BER of the MU-SIMO systems using ZF and ZF-SIC versus $M$ with mixed-precision arithmetic, $K=4$, $b=32$, SNR = 0 dB, and 16QAM.
Figure 13: Performance of different transceivers with mixed-precision arithmetic architecture.
VI-D5 Impact of Imperfect CSI
Channel estimation errors are inevitable in practice. Therefore, for the most practical and general case of imperfect CSI, we consider a transmission within the coherence interval $T$ and use $\tau$ symbols for pilots. The CSI for each antenna can be obtained using high-precision arithmetic. The power of pilot symbols is $\rho_{p}=\tau\rho$, where the MMSE estimate of $\bf H$ is given by $\hat{\bf H}$. Let $\Delta{\bf H}=\hat{\bf H}-{\bf H}$ be the channel estimate error matrix, which is independent of $\hat{\bf H}$. The entries of $\Delta{\bf H}$ are random variables with zero means and variances $\frac{1}{\rho_{p}+1}$ [1]. Following the simulation parameters in [1, 31], we set the coherence interval $T=196$, and pilot sequences of length $\tau=K$. Other parameters are the same as those in the perfect CSI scenario. As shown in Fig. 12a and 12b, the ergodic rates of SIMO and MU-SIMO systems under imperfect CSI are presented with different precision levels. It is clear that the proposed mixed-precision arithmetic architecture still performs well between full-precision and low-precision in the imperfect CSI scenario.
VI-D6 More Advanced Transceivers
More advanced precoding and receiver schemes are presented to demonstrate the superiority of the proposed mixed-precision arithmetic architecture, including linear transceiver methods, such as minimum mean-squared error (MMSE) [43], and non-linear transceiver methods, such as zero-forcing successive interference cancellation (ZF-SIC) [44], and weighted minimum mean-squared error (WMMSE) [45], as depicted in Fig. 13. It is clear that the proposed mixed-precision arithmetic architecture consistently performs well in terms of ergodic rate and bit error rate (BER).
@@ Appendix C Proof of Proposition 3
$\mathbb{E}\left\{\left\|\mathbf{r}\right\|_{2}^{2}\right\}$ $=\mathbb{E}\left\{\left\|\mathbf{A}^{H}\mathbf{y}\right\|_{2}^{2}\right\}=\mathbb{E}\left\{\left\|\sqrt{\rho}\mathbf{x}+\mathbf{A}^{H}\mathbf{n}\right\|_{2}^{2}\right\}$